OpenStudy (anonymous):

Find all solutions for this equation: 2cos^2(x)+sin(x)-2=0

OpenStudy (thomas5267):

TIp: \(\cos^2(x)+\sin^2(x)=1\)

OpenStudy (anonymous):

I'm sorry, I don't really understand how to use that. My teacher didn't explain it well.

OpenStudy (thomas5267):

\(\cos^2(x)=1-\sin^2(x)\) Is that more clear now?

OpenStudy (anonymous):

A little bit.

OpenStudy (thomas5267):

\[ \begin{align*} 2\cos^2(x)+\sin(x)-2&=0\\ 2(1-\sin^2(x))+\sin(x)-2&=0\\ 2-2\sin^2(x)+\sin(x)-2&=0\\ -2\sin^2(x)+\sin(x)&=0 \end{align*} \] More clear now?

OpenStudy (anonymous):

Yeah, so after doing that would I continue using pythagorean?

OpenStudy (thomas5267):

Nope. Just factor out \(\sin(x)\) and solve it.

OpenStudy (anonymous):

Ok, thank you!